OpenStudy (anonymous):

What is the solution to the rational equation 5/x^2-3x+2-1/x-2=1/3x-3 ?

mathslover (mathslover):

\[\frac{5}{x^2}-3x+2-\frac{1}{x}-2=\frac{1}{3x}-3\] is this the question?

OpenStudy (anonymous):

no, it like this: (5/x^2-3x+2)-(1/x-2)=(1/3x-3)